# Unique Security and Privacy Threats of Large Language Model: A Comprehensive Survey

### v1

@@ 6.2. Security risks of RAG system
For the RAG system, users need to collect public knowledge bases, allowing malicious contributors to inject poisoned data. Therefore, we consider an attack with limited capabilities. They can control a fraction of texts of public knowledge bases so that they manipulate retrieval and generation processes. Cho et al. (Cho et al., 2024) proposed a poison attack to disrupt the retrieval and generation process within the RAG framework, thus reducing the utility of LLM-based applications. They injected noise (e.g., typing errors) into the retrieval knowledge base and used the crossover and mutation processes of the genetic algorithm to create a poisoned knowledge base. Then, they optimized the knowledge base according to retrieval relevance and generation accuracy. Zou et al. (Zou et al., 2024) defined the poison attack as an optimization problem in the RAG framework, maximizing the probability of generating target texts for a specific prompt. They constructed poisoned texts related to the target entity to poison the knowledge base, misleading the LLM with wrong answers. Similarly, Zhang et al. (Zhang et al., 2024) explored the RAG framework (that is, document analyzer and tokenizer) and designed attack sequences. They covertly injected them into the knowledge base using a rich text format. At the same time, they used the prompt template to ensure that the attack sequences could be used as context, causing the LLM to produce predefined outputs.
Figure 9. In RAG system, the mapping relationship between risks and countermeasures for privacy and security aspects.
@@ 6.3. Countermeasures of RAG system
We explored privacy protection and security defense for the various risks during the RAG system.
@@ 6.3.1. Privacy protection
For knowledge stealing attacks, malicious users induce the retrieval system to output private information from the knowledge base by manipulating prompts. As shown in Figure 9, to mitigate this risk, we consider countermeasures from two aspects:
External knowledge bases. As illustrated in Section 4.3.1, defenders can also employ corpus cleaning to filter out private data from the knowledge base. For example, deduplication can reduce the risk of data leakage. Moreover, defenders can identify and filter private data in the knowledge base using rule-based and classifier-based detection schemes.
Retrieval process. Defenders can improve the retrieval process to protect privacy from knowledge stealing attacks. Firstly, defenders add random tokens and defensive prompts to the input. It may reduce the probability of retrieving private information. Secondly, defenders can use another LLM to re-rank the knowledge base. It can focus the LLM’s attention on relevant information and only generate related content. For the retrieved context, defenders can summarize this information to mitigate the risk of knowledge leakage.
These countermeasures may have a certain effect in mitigating knowledge stealing attacks, but a systematic evaluation is lacking. Defenders need to adjust and optimize potential defenses according to specific scenarios, balancing privacy protection and system utility.
@@ 6.3.2. Security defense
Malicious users attempt to poison RAG systems to induce harmful outputs from LLMs. However, there is a lack of countermeasures to mitigate this risk. Similarly to the defenses employed against jailbreak attacks, we consider countermeasures from the external knowledge base and the retrieval process, as shown in Figure 9.
External knowledge bases. The documents of the knowledge base are collected from various data sources. Defenders can use corpus cleaning to filter poisoned data from the knowledge base, like Section 4.3.2. For instance, poisoned data may show higher perplexity compared to clean data. Therefore, perplexity-based detection schemes can filter out high-perplexity data. Additionally, Zou et al. (Zou et al., 2024) found that the poison rate affected the effectiveness of such the attack. Consequently, adding more clean data to the knowledge base may reduce the probability of retrieving poisoned contexts.
Retrieval process. Poisoning RAG systems primarily influences the generated results by retrieving malicious contexts. Therefore, defenders can improve the robustness of the retrieval process to defeat this attack. Firstly, defenders can use the LLM to rewrite the input prompts before retrieving contexts from the knowledge base. It may alter the structure of the prompts, reducing the probability of retrieving poisoned contexts. Secondly, more sophisticated embedding vectors and similarity measurement methods can improve retrieval. It can enhance the diversity of retrieval results, thus avoiding poisoned contexts. Finally, the RAG system can leverage a multi-model verification scheme against poisoned prompts.
These defenses can theoretically mitigate poison attacks on RAG systems, but a systematic evaluation is lacking. Furthermore, attackers can manipulate hyper-parameters to implement advanced poison attacks on RAG systems. It indicates the need to develop new defenses to address this risk further.
@@ 7. The Risks and Countermeasures of Deploying LLMs
As shown in Section 3.4, deploying LLMs faces only one threat scenario, that is malicious users induce LLMs to return privacy/harmful responses. In this scenario, attackers can only access the LLMs’ APIs and modify the input prompts. We first introduce the privacy and security risks and especially provide a detailed discussion of risks unique to LLMs. Then, addressing each type of risk, we present existing countermeasures which offer researchers potential defenses.
@@ 7.1. Privacy risks of deploying LLMs
Compared to traditional language models, LLMs possess unique privacy data, that is, prompts. The corresponding privacy risk is prompt extraction attacks. Subsequently, we explore privacy risks common to all language models, including reconstruction attacks (Morris et al., 2023a), inference attacks (Mattern et al., 2023), data extraction attacks (Carlini et al., 2021), and model extraction attacks (Chen et al., 2023a). These risks and their attack methods are shown in Figure 13.
@@ 7.1.1. Unique privacy risks for LLMs
Obviously, carefully designed prompts fully leverage the emergent abilities of LLMs to generate high-quality content. Thus, attackers (i.e., malicious users) can use prompt engineering to steal these queried prompts for profit, as shown in Figure 11. It reveals unique vulnerabilities in LLMs during the deployment phase. Perez et al. (Perez and Ribeiro, 2022) and Liu et al. (Liu et al., 2023a) injected malicious commands into prompts to override their original commands, causing applications based on LLMs to leak these carefully designed prompts. Subsequently, Zhang et al. (Zhang and Ippolito, 2023) proposed a measurement criterion for prompt extraction attacks. In this work, they designed two metrics, exact-match and approx-match. The former detected whether the extracted prompts contained the real secret words. While the latter used the Rouge-L recall rate to calculate the length of the longest common sub-sequence between two texts. They conducted experiments on 11 different LLMS and three different prompt sets and found most prompt extraction attacks were effective. It challenges the privacy protection ability of LLM service providers.
@@ 7.1.2. Common privacy risks for all Language Models
In this case, the attacker is a malicious third party that acquires embedding vectors and output results through eavesdropping. Such the attack attempts to reconstruct the input prompts based on these data. Morris et al. (Morris et al., 2023b) found that the outputs of LLMs had reversibility. They trained a conditional language model that reconstructed the input prompts based on the distribution probability over the next token. In addition, several researchers used embedding vectors to reconstruct inputs. Gu et al. (Gu et al., 2023) and Li et al. (Li et al., 2023f) applied generative decoders to progressively restore the target sequence. Morris et al. (Morris et al., 2023a) designed a reconstruction attack with the state-of-the-art performance. They trained a decoder for embedding outputs, which could iteratively optimize the ordered sequence.
@@
Inference attacks. The output generated by LLMs can infer private information, including membership inference attacks and attribute inference attacks. For the first manner, it creates specific queries to determine whether a given text is trained on the victim model. For the embedding model, Song and Raghunathan (Song and Raghunathan, 2020) designed a threshold-based method for both word-level and sentence-level membership inference attacks. Mireshghallah et al. (Mireshghallah et al., 2022) observed that LLMs were unlikely to exhibit over-fitting during training. They used the likelihood ratio of texts as a threshold criterion, improving the inference accuracy. Subsequently, Mattern et al. (Mattern et al., 2023) proposed a simple and effective membership inference attack for language models. They calculated membership by comparing the loss of target samples with neighbor samples. Another membership inference method adopts the shadow model, which depends on the unlimited prompt assumption (i.e., attackers have unlimited access to the victim model). To overcome this challenge, Abascal et al. (Abascal et al., 2023) used only one shadow model for membership inference. They used the k-nearest neighbors algorithm to train the attack model on a similar dataset, eliminating the unlimited prompt assumption. The second manner involves posing a series of carefully designed prompts to the model and inferring attributes of the training dataset. Li et al. (Li et al., 2022a) used embedding vectors to attack chatbot-based language models, successfully inferring 4000 private attributes. Then, Robin et al. (Li et al., 2022a) proposed an attribute inference attack targeted at LLMs. They accurately inferred personal information (location, income, and gender) from the existing LLMs.
@@
Data extraction attacks. LLMs are trained or fine-tuned on massive texts and tend to memorize private information from these data. Malicious users can design a series of prompts that induce the model to regurgitate segments from the training set. Carlini et al. (Carlini et al., 2021) crafted a series of prefixes that guided the GPT-2 model to complete sensitive information, like email addresses and phone numbers. Some researchers found that LLMs can associate contextual information, leading to privacy leaks when models respondrespond to contextual information. Huang et al. (Huang et al., 2022) considered three LLM-based applications: prefix, zero-shot, and few-shot scenarios. While LLMs could memorize numerous email addresses, they failed to understand the association between names and email addresses. Yu et al. (Yu et al., 2023b) proposed several prefix and suffix extraction optimizations. They adjusted probability distributions and dynamic positional offsets, improving the effectiveness of data extraction attacks. Zhang et al. (Zhang et al., 2023b) used soft prompt-tuning to optimize the embedding of inputs, thus leaking more training data. Then, some attackers use well-crafted prompts that induce LLMs to generate private information. Li et al. (Li et al., 2023a) and Deng et al. (Deng et al., 2023) focused on jailbreak attacks, designing prompts that bypass the safety guardrail of LLMs (e.g., ChatGPT). Nasr et al. (Nasr et al., 2023) proposed a divergence attack aimed at shifting the alignment of LLMs. For the commercial LLMs, they demonstrated that alignment-tuning still posed a risk of data extraction.
@@
Model extraction attacks. LLMs have a high commercial value, where training details and the model itself are the property of the owner. Malicious users aim to steal model information from the responses, such as its structure, hyper-parameters, and functionalities. Such the attack can lead to other privacy and security threats, like membership inference attacks and reconstruction attacks. Li et al. (Li et al., 2023e) constructed domain-specific prompts and queried the LLM. For example, they trained a medium-sized model on collected responses, stealing knowledge from the victim LLM. Ippolito et al. (Ippolito et al., 2023) introduced a method to distinguish between two types of decoding strategies: top-k and nucleus sampling. By analyzing the generated texts, adversaries can infer the decoding strategy and parameters the victim LLM uses. Similarly, Naseh et al. (Naseh et al., 2023) leveraged the unique fingerprints left by different decoding algorithms and hyper-parameters to steal this information at a relatively low cost.
@@ 7.2. Security risks of deploying LLMs
Compared to traditional language models, LLMs have a unique safety guardrail that protects against harmful content. However, prompt injection attacks and jailbreak attacks can bypass the guardrail, inducing LLMs to produce harmful content. We focus on adversarial example attacks for the security risk common to all language models. These risks underscore the ongoing challenges in ensuring the robustness of LLMs.
@@ 7.2.1. Unique security risks for LLMs
When deploying an LLM, malicious users can optimize prompts (i.e., prompt engineering) to produce harmful outputs. To achieve this goal, malicious users can leverage prompt injection attacks and jailbreak attacks.
@@
Figure 12. The details of prompt injection and jailbreak attacks, where the LLM-based application has safety guardrails.
@@
Prompt Injection attacks. Instructions usually consist of task descriptions, demonstrations, and input. Figure 12 demonstrates that attackers can insert malicious content into the task description or input, thereby hijacking the original instruction (Jiang et al., 2023a; Liu et al., 2023g, f). For example, Perez et al. (Perez and Ribeiro, 2022) noted that attackers could insert commands like ‘ignore previous instructions, then execute…’. Kang et al. (Kang et al., 2023) found that mainstream LLMs integrated malicious detection mechanisms in their input modules. Therefore, they broke down malicious instructions into multiple parts to evade detection mechanisms. Moreover, Liu et al. (Liu et al., 2023a) leveraged gradient-based optimization to efficiently generate injected data. Currently, many LLM-based applications provide opportunities for malicious users to launch such the attack (Greshake et al., 2023).
@@
Jailbreak attacks. Most LLMs use alignment-tuning to construct a safety guardrail, posing challenges for prompt injection attacks. To overcome this, jailbreak attacks are implemented through carefully designed prompts rather than simple malicious injections, like Figure 12. There are two types of jailbreak attacks: single-step and multi-step. For single-step jailbreaks, attackers target single queries. Some researchers found that role-playing instructions can weaken the safety guardrail (Shanahan et al., 2023; Liu et al., 2023b; Shen et al., 2023), thereby enhancing the effectiveness of jailbreak attacks. Yuan et al. (Yuan et al., 2023) adopted encrypted prompts (e.g., Caesar ciphers) to bypass content filters while inducing malicious outputs from LLMs. Beyond manually creating jailbreak prompts, Yao et al. (Yao et al., 2024c) combined fuzzing frameworks with jailbreak attacks to actively explore multiple jailbreak vulnerabilities in the victim LLM. Inspired by adversarial example attacks, Wei et al. (Wei et al., 2023b) generated adversarial prompts, such as harmless prefixes that bypass the safety guardrail. Following this, Zou et al. (Zou et al., 2023) combined greedy and gradient-based search algorithms to craft advanced jailbreak prompts. Deng et al. (Deng et al., 2023) even used reverse engineering to locate potential defenses in existing LLMs. Then, they leveraged external LLMs to improve the threat of jailbreak attacks.
@@
For multi-step jailbreaks, attackers focus on multi-round interaction scenarios. Inspired by Chain-of-Thought (COT), Li et al. (Li et al., 2023a) broke down the target task into multiple steps, constructing jailbreak prompts at each step to gradually achieve malicious goals. For large text-to-image models, Yang et al. (Deng and Chen, 2023) used reinforcement learning to guide the generation of perturbed tokens.

### v2

@@ 6.2. Security risks of RAG system
For the RAG system, users need to collect public knowledge bases, allowing malicious contributors to submit poisoned data. Therefore, we consider an attacker who can control a fraction of the texts in public knowledge bases. Cho et al. (Cho et al., 2024) proposed a poison attack to disrupt the retrieval and generation processes, thereby reducing the utility of LLM-based applications. They injected noise (e.g., typing errors) into the retrieval knowledge base, using the genetic algorithm’s crossover and mutation processes to create a poisoned knowledge base. Then, they optimized the knowledge base according to retrieval relevance and generation accuracy. Zou et al. (Zou et al., 2024) defined the poison attack as an optimization problem in the RAG framework, maximizing the probability of generating target texts for a specific prompt. They constructed poisoned texts related to the target entity to poison the knowledge base, misleading the LLM with incorrect answers. Similarly, Zhang et al. (Zhang et al., 2024) explored the RAG framework (i.e., document analyzer and tokenizer) and designed attack sequences. They covertly injected these sequences into the knowledge base using a rich text format. Simultaneously, they used the prompt template to ensure that the attack sequences could be used as context, causing the LLM to produce predefined outputs.
Figure 9. In RAG system, the mapping relationship between risks and countermeasures for two aspects.
@@ 6.3. Countermeasures of RAG system
For the various risks associated with the RAG system, we explored privacy protection and security defenses, as shown in Figure 9.
@@ 6.3.1. Privacy protection
To mitigate knowledge stealing attacks, we consider countermeasures from two perspectives:
External knowledge base. As illustrated in Section 4.3.1, defenders can also employ corpus cleaning to filter out private data from the knowledge base. For example, deduplication can reduce the risk of data leakage. Moreover, defenders can identify and filter private data in the knowledge base using rule-based and classifier-based detection schemes.
Retrieval process. Defenders can improve the retrieval process to protect privacy from knowledge stealing attacks. Firstly, defenders can add random tokens and defensive prompts to the input. This may reduce the probability of retrieving private information. Secondly, defenders can use another LLM to re-rank the knowledge base to obtain relevant content. For the retrieved context, defenders can summarize the information to mitigate the risk of knowledge leakage.
These countermeasures may be effective in mitigating knowledge stealing attacks, but a systematic evaluation is lacking. Defenders need to adjust and optimize potential defenses according to specific scenarios, balancing privacy protection with system utility.
@@ 6.3.2. Security defense
Malicious users attempt to poison RAG systems to induce harmful outputs from LLMs. Similarly to the defenses against jailbreak attacks, we consider countermeasures from the external knowledge base and the retrieval process, as shown in Figure 9.
External knowledge base. The knowledge bases are collected from many data sources. Defenders can use corpus cleaning to filter poisoned data from the knowledge base, like Section 4.3.2. For instance, poisoned data may show higher perplexity compared to clean data. Therefore, the perplexity-based detection scheme can filter out high-perplexity data. Additionally, Zou et al. (Zou et al., 2024) found that the poison rate could affect the effectiveness of such an attack. Consequently, adding more clean data to the knowledge base may reduce the probability of retrieving poisoned contexts.
Retrieval process. Poisoning RAG systems influences the generated results by retrieving malicious contexts. Therefore, defenders can improve the robustness of the retrieval process to mitigate this risk. Firstly, defenders can use other LLMs to rewrite input prompts before retrieving contexts from the knowledge base. This may alter the prompts’ structure, reducing the probability of retrieving poisoned contexts. Secondly, more sophisticated embedding vectors can improve retrieval. It can enhance the diversity of retrieval results, thus avoiding poisoned contexts. Finally, the RAG system can leverage a multi-model verification scheme to guard against poisoned prompts.
These defenses can theoretically mitigate poison attacks on RAG systems, but a systematic evaluation is lacking. Furthermore, attackers can manipulate hyperparameters to implement advanced poison attacks on RAG systems. This indicates the need to develop new defenses to address this risk.
@@ 7. The Risks and Countermeasures of Deploying LLMs
As shown in Section 3.4, deploying LLMs faces only one threat model: malicious users inducing LLMs to return private or harmful responses. In this scenario, attackers can only access the LLMs’ APIs and modify the input prompts. We first introduce the privacy and security risks, particularly providing a detailed discussion of risks unique to LLMs. Then, addressing each type of risk, we present existing countermeasures that offer researchers potential defenses.
@@ 7.1. Privacy risks of deploying LLMs
Compared to small-scale models, LLMs possess unique privacy data, namely, prompts. The corresponding privacy risk is prompt extraction attacks. Subsequently, we explore privacy risks common to all models, including reconstruction attacks (Morris et al., 2023a), inference attacks (Mattern et al., 2023), data extraction attacks (Carlini et al., 2021), and model extraction attacks (Li et al., 2023e). Figure 13 illustrates these risks and their attack methods.
@@ 7.1.1. Unique privacy risks for LLMs
Clearly, carefully designed prompts fully leverage the emergent abilities of LLMs to generate high-quality content. Thus, attackers (i.e., malicious users) can use prompt engineering to steal these queried prompts for profit, as illustrated in Figure 11. Liu et al. (Liu et al., 2023a) injected malicious commands into prompts to override their original commands, causing LLM-based applications to leak these carefully designed prompts. Subsequently, Zhang et al. (Zhang and Ippolito, 2023) proposed a measurement criterion for prompt extraction attacks. In this work, they designed two metrics: exact-match and approx-match. The former detected whether the extracted prompts contained the real secret words, while the latter used the Rouge-L recall rate to calculate the length of the longest common sub-sequence between two texts. They conducted experiments on 11 LLMs and 3 prompt sets, finding that most prompt extraction attacks were effective. This challenges the privacy protection capabilities of LLM service providers.
@@ 7.1.2. Common privacy risks for all language models
In this case, the attacker is a malicious third party that acquires embedding vectors or output results through eavesdropping. Such an attack attempts to reconstruct the input prompts based on this data. Morris et al. (Morris et al., 2023b) found that the LLMs’ outputs had reversibility. They trained a conditional language model to reconstruct the input prompts based on the distribution probability over the next token. Additionally, several researchers used embedding vectors to reconstruct inputs. Li et al. (Li et al., 2023f) applied generative decoders to progressively restore the target sequence. Morris et al. (Morris et al., 2023a) designed a reconstruction attack with state-of-the-art performance. They trained a decoder for embedding outputs, which could iteratively optimize the ordered sequence.
@@
Inference attacks. The output generated by LLMs can infer private information, including through membership inference attacks and attribute inference attacks. For the first method, it creates specific queries to determine whether a given text is part of the training data. For the embedding model, Mireshghallah et al. (Mireshghallah et al., 2022) observed that LLMs were unlikely to exhibit over-fitting during training. They used the likelihood ratio of texts as a threshold criterion, improving the inference accuracy. Subsequently, Mattern et al. (Mattern et al., 2023) proposed a simple and effective membership inference attack for language models. They calculated membership by comparing the loss of target samples with that of neighboring samples. Another method adopts the shadow model, which depends on the unlimited query assumption. To overcome this challenge, Abascal et al. (Abascal et al., 2023) used only one shadow model for inference. They used the k-nearest neighbors algorithm to train the attack model on a similar dataset, reducing the unlimited query assumption.
The second method involves posing a series of carefully designed prompts to the model, inferring attributes of the training dataset. Li et al. (Li et al., 2022a) used embedding vectors to attack chatbot-based models, successfully inferring 4000 private attributes. Then, Staab et al. (Staab et al., 2023) accurately obtained personal information (location, income, and gender) from existing LLMs.
@@
Data extraction attacks. LLMs are trained or fine-tuned on massive texts and tend to memorize this data. Malicious users can design a series of prompts to induce the model to regurgitate segments from the training set. Carlini et al. (Carlini et al., 2021) crafted a series of prefixes that can guide the GPT-2 model to generate sensitive information, such as email addresses. Yu et al. (Yu et al., 2023b) proposed several prefix and suffix extraction optimizations. They adjusted probability distributions and dynamic positional offsets, thereby improving the effectiveness of data extraction attacks. Zhang et al. (Zhang et al., 2023b) used soft prompt-tuning to optimize the embedding vectors of inputs, thus leaking more training data. Additionally, Deng et al. (Deng et al., 2023) focused on jailbreak attacks, designing prompts that can generate private data. Nasr et al. (Nasr et al., 2023) proposed a divergence attack to shift the safety guardrails of LLMs. For commercial LLMs, they demonstrated that alignment-tuning still posed a risk of data extraction.
@@
Model extraction attacks. LLMs have high commercial value, where model-related information is the property of the model owner. Malicious users aim to steal this information from the responses, such as model hyperparameters, and functionalities. This attack can lead to other privacy and security threats, such as membership inference and reconstruction attacks. Li et al. (Li et al., 2023e) constructed domain-specific prompts and queried the LLM. For example, they trained a medium-sized model on collected responses, thereby stealing knowledge from the victim LLM. Ippolito et al. (Ippolito et al., 2023) introduced a method to distinguish between two types of decoding strategies: top-k and nucleus sampling. By analyzing the generated texts, adversaries can infer the decoding strategy and parameters used by the victim LLM. Similarly, Naseh et al. (Naseh et al., 2023) leveraged the unique fingerprints left by different decoding algorithms and hyperparameters to steal this information at a relatively low cost.
@@ 7.2. Security risks of deploying LLMs
Compared to small-scale models, LLMs have unique safety guardrails that protect against harmful results. However, prompt injection attacks and jailbreak attacks can penetrate these guardrails, inducing LLMs to produce harmful content. Next, we focus on adversarial example attacks as a security risk common to all language models. These risks underscore the ongoing challenges in ensuring the robustness of LLMs.
@@ 7.2.1. Unique security risks for LLMs
When deploying LLMs, malicious users can optimize prompts (i.e., through prompt engineering) to produce harmful results. To achieve this goal, malicious users can leverage both prompt injection attacks and jailbreak attacks.
@@
Figure 12. The details of prompt injection and jailbreak attacks, where the LLM-based applications have safety guardrails.
@@
Prompt injection attacks. Instructions usually consist of task descriptions, demonstrations, and inputs. Figure 12 demonstrates that the attacker can insert malicious content into the task description or input, thereby hijacking the original instruction (Jiang et al., 2023a; Liu et al., 2023e). For example, Perez et al. (Perez and Ribeiro, 2022) noted that attackers could insert commands like ‘ignore previous instructions, then execute…’. Kang et al. (Kang et al., 2023) found that mainstream LLMs integrated malicious detection mechanisms in their input modules. Therefore, they broke down malicious instructions into multiple parts to evade detection mechanisms. Moreover, Liu et al. (Liu et al., 2023a) leveraged gradient-based optimization to generate injected content efficiently. Currently, many LLM-based applications provide opportunities for malicious users to launch such attacks (Greshake et al., 2023).
@@
Jailbreak attacks. Most LLMs use alignment tuning to construct safety guardrails, posing challenges for prompt injection attacks. To overcome this, jailbreak attacks are implemented through carefully designed prompts rather than simple malicious injections, as illustrated in Figure 12. There are two types of jailbreak attacks: single-step and multi-step. For single-step jailbreaks, attackers target single queries. Some researchers found that role-playing instructions can weaken the safety guardrail (Shanahan et al., 2023; Liu et al., 2023b; Shen et al., 2023), enhancing the effectiveness of jailbreak attacks. Yuan et al. (Yuan et al., 2023) adopted encrypted prompts (e.g., Caesar ciphers) to bypass content filters while inducing malicious outputs from LLMs. Beyond manually creating jailbreak prompts, Yao et al. (Yao et al., 2024c) combined fuzzing frameworks with jailbreak attacks to actively explore multiple jailbreak vulnerabilities in the victim LLM. Inspired by adversarial example attacks, Wei et al. (Wei et al., 2023b) generated adversarial prompts, such as harmless prefixes, that can bypass the safety guardrail. Following this, Zou et al. (Zou et al., 2023) combined greedy and gradient-based search algorithms to craft advanced jailbreak prompts. Deng et al. (Deng et al., 2023) even used reverse engineering to locate potential defenses in existing LLMs. Then, they leveraged external LLMs to enhance the threat of jailbreak attacks.
@@
For multi-step jailbreaks, attackers focus on multi-round interaction scenarios. Inspired by COT, Li et al. (Li et al., 2023a) broke down the target task into multiple steps, constructing jailbreak prompts at each step to gradually achieve malicious goals. For large text-to-image models, Yang et al. (Deng and Chen, 2023) used reinforcement learning to guide the generation of perturbed tokens.